OpenStudy (anonymous):

Medal will be awarded for Algebra 1

OpenStudy (anonymous):

I really need help on these!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Is this a rolling dice question?

OpenStudy (anonymous):

yes it is hold on

OpenStudy (anonymous):

A number cube with the numbers 1 through 6 is rolled. Find the given probability.

OpenStudy (anonymous):

srry idk what is going on with my computer

OpenStudy (anonymous):

there

OpenStudy (anonymous):

Okay Sample Space = 1, 2, 3, 4, 5, and 6 1. P(Less than 2) = 1/6 2. P(Greater than/Equal to 3) = 3,4,5,6 = 4/6 = 2/3 3. P(Complement of 4) = 1,2,3,5,6 = 5/6 4. 1:4

OpenStudy (anonymous):

@Dharmaputra_993 isn't #4 1:5 because there are 5 possible answers?

OpenStudy (anonymous):

But for 1 to be right, other 4 must be wrong =)

OpenStudy (anonymous):

Oh okay now I understand! Thanks!